OpenStudy (anonymous):

What would be the next number in pattern 1/324, 1/108, 1/36, 1/12 answers are 1/3, 1/5, 1/4, 1

OpenStudy (kc_kennylau):

Try finding the common ratio of this sequence :)

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

it's actually 1/3

OpenStudy (anonymous):

@DVReed96 you are wrong.

OpenStudy (kc_kennylau):

no he's referring to the answer instead of to my question

OpenStudy (anonymous):

yeah the answer is 1/3 not 1/4 because 1/3 times 1/12 equals 1/36

OpenStudy (kc_kennylau):

No the answer is 1/4 because the common ratio is 3

OpenStudy (kc_kennylau):

1/324 * 3 = 1/108 1/108 * 3 = 1/36 1/36 * 3 = 1/12